The perimeter of a semicircle is doubled when the radius is increased by 3. Find the radius of the semicircle.
Answer by Alan3354(69443)

The perimeter of a semicircle is doubled when the radius is increased by 3. Find the radius of the semicircle.
p = pi*r + 2r
= r*(pi+2)
----------------
2p = (r+3)*(pi+2)
p = r*(pi+2)
----------------- Subtract
p = 3*(pi+2) = r*(pi+2)
r = 3